A rectangular block of metal of 50 mm x 25 mm cross-section and 125 mm length carries a tensile load of 100 kN along its length, a compressive load of 1.0 MN on its 50 x 125 mm faces and a tensile load of 400 kN on its 25 x 125 mm faces. If E = 208 GN/m² and Poisson’s ratio = 0.3, Find (a) change in volume of bar (b) increase required in 1.0 MN load to produce no change in volume.
1
Expert's answer
2020-08-17T09:23:26-0400

As per the question,

for rectangular block,

length=125mm

breadth=50mm

height=25mm


Load along its length=100 KN

Load along its breadth= -1 MN=-1"\times10^3" KN( since It is compressive load)

Load along its height=400 KN

Young's modulus=208GN/m"^2" =208"\frac{\times 10^6}{10^6}" KN/mm"^2" =208KN/mm"^2"

poission's ratio=0.3


Stress in the x direction "\sigma_x" ="\frac{load in x direction}{breadth\times height}"

="\frac{100}{25\times 50}"

=80 N/mm"^2"

Stress in the y direction "\sigma_y" ="\frac{load in y direction}{lenght\times breadth}"

= -"\frac{1000}{50\times125}"

=-100 N/mm"^2"

Stress in the z direction "\sigma_z" ="\frac{load in z direction}{length\times height}"

="\frac{100}{125\times50}"

=128 N/mm"^2"

Let the volume of block be V,

then V="length\times breadth\times height"

="125\times 50\times 25"

=156250 mm"^3"

Now the change in volume to original volume is given by,

"\frac{dV}{V}= \frac{(\sigma_x -\sigma_y+\sigma_z)(1-2u)}{E}"

="\frac{(80-100+128)(1-2\times 0.3)}{208\times 10^3}"

=0.0002076

Now the change in volume,

dV =0.0002076"\times" V

=0.0002076"\times" 156250

=32.4375 mm"^3"

Hence the change in volume is 32.4375mm"^3"




(B) To keep the volume of the Block zero, we have to make the net load applied to zero.

This can be done by making the compressive 1 MN load to 0.5 MN.